Find to three significant digits the charge and the mass of the following particles. Suggestion: Begin by looking up the mass of a neutral atom on the periodic table of the elements in Appendix C (a) an ionized hydrogen atom, represented as H charge 0 1.602e-19x Check the number of significant figures. C mass 1.67e-27 kg (b) a singly ionized sodium atom, Na charge 492 1.6020-19x Check the number of significant figures. C 401.67e-27x Your response differs significantly from the correct answer. Rework your solution from the beginning and check each step carefully. kg mass (c) a chloride ion CI 41.602e-19X charge The response you submitted has the wrong sign. C 4 9.11e-31X mass Your response differs significantly from the correct answer. Rework your solution from the beginning and check each step carefully. kg (d) a doubly ionized calcium atom, CaCa charge 3.20e-19 C mass 403.340-27X Your response differs significantly from the correct answer. Rework your solution from the beginning and check each step carefully. kg (e) the center of an ammonia molecule, modeled as an N3- ion 44.01e-19X charge The response you submitted has the wrong sign. C 402.73e-31 X mass Your response differs significantly from the correct answer. Rework your solution from the beginning and check each step carefully. kg (f) quadruply ionized nitrogen atoms, N4+, found in plasma in a hot star charge126.41e-19 C 6.68e-27 X mass Your response differs from the correct answer by more than 10%. Double check your calculations. kg (9) the nucleus of a nitrogen atom charge 40 1.12e-19x Your response is off by a multiple of ten. C 402 3.255e-15 X Your response differs significantly from the correct answer. Rework your solution from the beginning and check each step carefully. kg mass (h) the molecular ion H2o 40 1.6e-19x The response you submitted has the wrong sign. C 4 9.11e-31 X Your response differs significantly from the correct answer. Rework your solution from the beginning and check each step carefully. kg charge mass
Answer #1

magnitude of charge on electron/proton = 1.60*10^-19 C

mass of a nucleon atom = 1.67*10^-27 kg

(a) given atom = H+

charge on H+ = 1*(charge on proton)

charge on H+ = 1.60*10^-19 C

there are 1 neutron in H+ ,

So, mass of H+ = 1.67*10^-27 kg

(b.) Na+

charge on Na+ = 1*(charge on proton)

charge on Na+ = 1.60*10^-19 C

Since there are 23 nucleon in Na+

So, mass of Na+ = 23*1.67*10^27 kg

mass of Na+ = 3.84*10^-26 kg

(c.) Cl-

charge on Cl- = 1*(charge on electron)

charge on Cl- = -1.60*10^-19 C

Since there are 35 nucleon in Cl-

So, mass of Cl- = 35*1.67*10^27 kg

mass of Cl- = 5.85*10^-26 kg

(d.) Ca++

charge on Ca++ = 2*(charge on proton)

charge on Ca++ = 3.20*10^-19 C

Since there are 40 nucleon in Ca++

So, mass of Ca++ = 40*1.67*10^27 kg

mass of Ca++ = 6.68*10^-26 kg

(e.) N3-

charge on N3- = 3*(charge on electron)

charge on N3- = -4.80*10^-19 C

Since there are 14 nucleon in N3-

So, mass of N3- = 14*1.67*10^27 kg

mass of N3- = 2.34*10^-26 kg

(f.) N4+

charge on N4+ = 4*(charge on proton)

charge on N4+ = -6.40*10^-19 C

Since there are 14 nucleon in N4+

So, mass of N4+ = 14*1.67*10^27 kg

mass of N4+ = 2.34*10^-26 kg

(g.) N

charge on N = 0*(charge on electron)

charge on N = 0 C

Since there are 14 nucleon in N

So, mass of N = 14*1.67*10^27 kg

mass of N = 2.34*10^-26 kg

(h.) H2O-

charge on H2O- = 1*(charge on electron)

charge on H2O- = -1.60*10^-19 C

Since there are 18 nucleon in H2O-

So, mass of H2O- = 18*1.67*10^27 kg

mass of H2O- = 3.01*10^-26 kg

Please upvote.